OpenStudy (rsadhvika):

The primes 37,67,73,79,... are of the form p = 36ab + 6a - 6b + 1, with a >= 1, b >=1 Show that no pair of twin primes can contain a prime of this form.

OpenStudy (ikram002p):

Let p,q be twin primes. The difference would be 36(ab-a2b2)+6(a-a1)-6(b-b2)

OpenStudy (ikram002p):

The least difference is 6

OpenStudy (ikram002p):

No wait lol

OpenStudy (ikram002p):

Ok here again.. I'll fix my mistake

OpenStudy (epoweritheta):

Looks bit tricky ...but easy :) Ok let us assume that we have a twin prime of this form call that p, then p =36ab + 6a - 6b + 1 (for some a and b) then since we assumed this guy p is a twin prime then p-2 or p+2 is prime . Now , we are required to prove that both these will not be prime ie if p is a prime (of that form) then p+2 & p-2 are not-prime Lets start 1) proving p+2 is composite(not-prime) whenever this guy p is prime . seen easily by adding 2 see : p =36ab + 6a - 6b + 1 +2 =p =36ab + 6a - 6b + 3 which has a factor of 3 . [end of 1] 2) proving p-2 requires a bit of manipulation techniques . see if p =36ab + 6a - 6b +1 then p-2=p =36ab + 6a - 6b - 1 which can be factorized into (6b+1)*(6a-1) since p=6a(6b+1)-(6b+1) =(6a-1)*(6b+1) [end of 2] QED

OpenStudy (ikram002p):

p and q are twins assume p is of the form q= 36ab + 6a - 6b +3 which is never prime xD

OpenStudy (epoweritheta):

^ i agree :)

OpenStudy (ikram002p):

well no need to other way lol @epoweritheta already did it..

OpenStudy (epoweritheta):

why what if in the twin primes pairs (ordered) <p,q> . you only proved p cannot be of the form 36ab + 6a - 6b + 1 . but what if q is of such form ??? you just proved q+2 is not prime but you need to prove p is not a prime .

OpenStudy (ikram002p):

=) yeah i was gonna write the rest but u already did :P

OpenStudy (epoweritheta):

:)